i"m sorry i missed another question
* the bears set out to pick berries for one of carol's famous pies. when they stopped picking:
*one bear sampled a berry and then ate 1/2 of what was left.
*another bear sampled a berry and ate 1/2 of whatwas left.
maria took the bucket and saw only twelve berries left. how many berries did they pick in all

Answer by scott8148(6628)

let x=berries picked

one bear sampled a berry and then ate 1/2 of what was left ... (x-1)(1/2)

another bear sampled a berry and then ate 1/2 of what was left ... ((x-1)(1/2)-1)(1/2)

maria took the bucket and saw only twelve berries left ... =12

((x-1)(1/2)-1)(1/2)=12 ... ((x-1)(1/2)-1)=24 ... (x-1)(1/2)=25 ... x-1=50 ... x=51